OpenStudy (anonymous):

calculus help please find the lines that are (a) tangent and (b) normal to the curve y=sqrt(x) at x=4

OpenStudy (anonymous):

y=x^1/2

OpenStudy (anonymous):

find derivative

OpenStudy (anonymous):

1/2sqrtx put x=4 1/(2)(2) slope=1/4

OpenStudy (anonymous):

of tangent at x=4

OpenStudy (anonymous):

ok so the slope is 1/4 so then the equation of the tangent would be y-2=(1/4)(x-4) which is really y=(1/4)x +1??

OpenStudy (anonymous):

and then since the equation of the normal has an opposite and reciprocal slope would the equation of the normal be y=-4x+16??

OpenStudy (anonymous):

y=2 you are right

OpenStudy (anonymous):

for normal line take - inverse of slope of tangen

OpenStudy (anonymous):

ok..so now would this be correct? EQ tangent = y=(1/4)x +1 EQ normal = y=-4x+16

OpenStudy (anonymous):

slope of normall =-4 pass through 4,2 y-2=-4(x-2) y=-4x+8+2 y=-4x+10

OpenStudy (anonymous):

ok that makes perfect sense. i guess i did my quick algebra incorrectly. thank you so much you have been a great help

OpenStudy (anonymous):

welcome